# MiniAppBench: Evaluating the Shift from Text to Interactive HTML Responses in LLM-Powered Assistants Zuhao Zhang1,2*, Chengyue Yu1*, Yuante Li3, Chenyi Zhuang1†, Linjian Mo1, Shuai Li2

### 10 Mar 2026 v1

@@ Abstract
[…]
MiniAppBench: Evaluating the Shift from Text to
Interactive HTML Responses in LLM-Powered Assistants
Zuhao Zhang1,2*, Chengyue Yu1*, Yuante Li3, Chenyi Zhuang1†, Linjian Mo1, Shuai Li2
1Inclusion AI, Ant Group 2Shanghai Jiao Tong University 3Carnegie Mellon University
@@ 1 Introduction
This transformation facilitates a paradigm shift in human-LLM interaction (as illustrated in Figure 1), moving from static text-only responses to rich, code-based engagements. Users now expect LLMs to produce interactive visualizations or functional applications that embody real-world logic. Consequently, to ensure these interactions feel natural and seamless, the model must actively capture and construct implicit assumptions or principles, such as “an object in free fall follows Newton’s laws" or “a week has seven days", which, while often taken for granted in human communication, are essential for valid execution. Real-world cases are shown in Figure 2.
@@
In practice, achieving these properties is non-trivial. As shown in Figure 2, an artifact may be syntactically valid and successfully executable, but still fail to support high-fidelity, non-fragmented interaction aligned with real user reasoning. To bridge this gap, we introduce MiniAppBench, the first benchmark designed specifically to evaluate the ability of LLMs to generate MiniApps. The comparison with other benchmarks is provided in the Appendix A. MiniAppBench is constructed through a rigorous multi-stage pipeline that distills tens of millions of real-world user queries into a balanced set of principle-driven, interaction-intensive tasks.
@@ 5.1 Settings
Baseline models are selected to ensure breadth, currency, and reproducibility, considering: (1) multiple model families (Claude (Anthropic, 2025a, b), Gemini (Google DeepMind, 2025a, b), GLM (Zeng et al., 2025), GPT (OpenAI, 2025b, a), Grok (xAI, 2025), Hunyuan (Team et al., 2025b), Kimi (Team et al., 2025a), Mimo (Xiao et al., 2026), MiniMax (Chen et al., 2025), and Qwen3 (Yang et al., 2025)); (2) a range of scales (from lightweight to flagship); and (3) relatively recent and representative versions within each family. For evluation, we select Gemini-3-pro as the evaluation-model driving agent due to its strong agreement with human judgments.
@@ C.1.3 Positive/Negative Labeling
We convert the three-dimensional evaluation scores into a binary label for downstream analysis. A MiniApp is marked as positive (successful) if all three dimension scores exceed a predefined threshold, i.e.,
(we use in the main setting). Otherwise, it is labeled as negative (failed). This conservative rule ensures that a sample is counted as successful only when it simultaneously satisfies the user intention, static correctness, and dynamic interaction requirements.

### 16 Mar 2026 v2

# MiniAppBench: Evaluating the Shift from Text to Interactive HTML Responses in LLM-Powered Assistants Zuhao Zhang 1 2 * Chengyue Yu 1 * Yuante Li 3 Chenyi Zhuang 1 † Linjian Mo 1 Shuai Li 2

@@ 1 Introduction
This transformation facilitates a paradigm shift in human-LLM interaction (as illustrated in Figure 1), moving from static text-only responses to rich, code-based engagements. Users now expect LLMs to produce interactive visualizations or functional applications that embody real-world logic. Consequently, to ensure these interactions feel natural and seamless, the model must actively capture and construct implicit assumptions or principles, such as “an object in free fall follows Newton’s laws” or “a week has seven days”, which, while often taken for granted in human communication, are essential for valid execution. Real-world cases are shown in Figure 2.
@@
In practice, achieving these properties is non-trivial. As shown in Figure 2, an artifact may be syntactically valid and successfully executable, but still fail to support high-fidelity, non-fragmented interaction aligned with real user reasoning.
To bridge this gap, we introduce MiniAppBench, the first benchmark designed specifically to evaluate the ability of LLMs to generate MiniApps. The comparison with other benchmarks is provided in the Appendix A. MiniAppBench is constructed through a rigorous multi-stage pipeline that distills tens of millions of real-world user queries into a balanced set of principle-driven, interaction-intensive tasks.
@@ 5.1 Settings
Baseline models are selected to ensure breadth, currency, and reproducibility, considering: (1) multiple model families (Claude (Anthropic, 2025a; b), Gemini (Google DeepMind, 2025a; b), GLM (Zeng et al., 2025), GPT (OpenAI, 2025b; a), Grok (xAI, 2025), Hunyuan (Team et al., 2025b), Kimi (Team et al., 2025a), Mimo (Xiao et al., 2026), MiniMax (Chen et al., 2025), and Qwen3 (Yang et al., 2025)); (2) a range of scales (from lightweight to flagship); and (3) relatively recent and representative versions within each family. For evluation, we select Gemini-3-Pro-Preview as the evaluation-model driving agent due to its strong agreement with human judgments.
@@ C.1.3 Positive/Negative Labeling
We convert the three-dimensional evaluation scores into a binary label for downstream analysis. A MiniApp is marked as positive (successful) if all three dimension scores exceed a predefined threshold, i.e., (we use in the main setting). Otherwise, it is labeled as negative (failed). This conservative rule ensures that a sample is counted as successful only when it simultaneously satisfies the user intention, static correctness, and dynamic interaction requirements.